OpenStudy (anonymous):

State how many imaginary and real zeros the function has. x^3 + 5x^2 + x + 5

OpenStudy (jdoe0001):

look at the signs changes, to use Descartes Law of Signs x^3 + 5x^2 + x + 5 ^ ^ ^ ^ + + + +

OpenStudy (jdoe0001):

that'd give you the "real positive" zeros now for for the "real negative" zeros let's get f(-x)

OpenStudy (jdoe0001):

f(-x) -x^3 + 5x^2 - x + 5 ^ ^ ^ ^ - + - +

OpenStudy (anonymous):

what do i have to do though to get these zeros?

OpenStudy (jdoe0001):

x^3 + 5x^2 + x + 5 ^ ^ ^ ^ + + + + no no no so, no real positive zeros

OpenStudy (whpalmer4):

We can use Descartes' Rule of Signs to find this out. Take the polynomial written in standard form (decreasing powers of the variable) and count the number of sign changes encountered when scanning left to right: ++++ are the signs of the coefficients of f(x), so there are 0 sign changes. That means we have 0 positive, real roots. Now rewrite the polynomial as f(-x). In effect, just change the signs of the odd-powered terms: -x^3+5x^2-x+5 Count again in the same fashion -+-+ 3 sign changes, so we have 3 negative real roots, possibly reduced by a multiple of 2. Complex roots always come in conjugate pairs in polynomials with only real coefficients (no \(i\)'s lurking in the coefficients). We also know that we always have exactly as many roots as the highest exponent of the variable. The difference between the count of real roots from the first two steps and the total is made up by complex roots, again, always in conjugate pairs. So, we have 0 positive real roots, and either 3 negative, real roots or 1 negative real root + 2 complex roots.

OpenStudy (jdoe0001):

descartes law of signs, is to see how many times the "sign" changes from term to term

OpenStudy (whpalmer4):

Actually finding the values of the roots or zeros is an entirely different matter!

OpenStudy (anonymous):

so there are 3 real zeros and 0 imaginary zeros right?

OpenStudy (whpalmer4):

You can use the Rational Root Theorem to make some intelligent guesses as to possible roots. In the case of this polynomial, 1, -1, 5 and -5 would be trial roots to evaluate. If the polynomial evaluates to 0 at x = 1, -1, 5, or -5, then that value of x is a root. (aka zero).

OpenStudy (jdoe0001):

f(-x) -x^3 + 5x^2 - x + 5 ^ ^ ^ ^ - + - + yes yes yes it changed 3 times from negative to positive then from positive to negative then again from negative to positive 3 times means, it has either 3 OR 3-2 real negative zeros, so, 3 or 1

OpenStudy (whpalmer4):

No, actually this polynomial has 1 negative real root and 2 complex roots.

OpenStudy (anonymous):

so 1 real zero and 2 imaginary zeros?

OpenStudy (jdoe0001):

right it can have EITHER 3 negative roots, in which case that'd leave room for no complex ones OR if it has 1 negative one, then you'd have room for 2 complex ones

OpenStudy (anonymous):

so which is the answer the 1 real zero and 2 imaginary zeros ; or 3 real zeros and 0 imaginary zeros

OpenStudy (jdoe0001):

that IS the answer :), both, either this or that

OpenStudy (jdoe0001):

yes, is kinda ambiguous, but descartes law of signs IS ambiguous, but it gives you a good ballpark figure on what you'd get :)

OpenStudy (jdoe0001):

look at it this way, is better than 0

OpenStudy (whpalmer4):

If you try out those possible roots I suggested (courtesy of the RRT): \[(1)^3+5(1)^2+(1)+5 = 12\] \[(-1)^3+5(-1)^2+(-1)+5=8\] \[(5)^3+5(5)^2+(5)+5=260\] \[(-5)^3+5(-5)^2+(-5)+5=0\] So x = -5 is a root. Thanks to the factor theorem, we can divide the polynomial by (x-(-5)) and apply the whole procedure again to get the remaining roots. \[\frac{x^3+5x^2+x+5}{x+5} = x^2+1\] Now you can find the roots of that, I'm sure, and you'll see that they are imaginary.

OpenStudy (anonymous):

well it is multiple choice so which one do i pick?

OpenStudy (jdoe0001):

what are your choices?

OpenStudy (whpalmer4):

Descartes gives you the possible combinations. You do have other tools which will clarify the situation, such as what I just did.

OpenStudy (anonymous):

0 imaginary; 3 real 1 imaginary; 2 real 3 imaginary; 0 real 2 imaginary; 1 real

OpenStudy (jdoe0001):

the first and the last choice are good, so I guess if you had to pick one, could be either because the function will yield either based on the descartes law of signs

OpenStudy (whpalmer4):

the middle two choices are impossible — with a real coefficient polynomial, the imaginary or complex roots MUST come in pairs.

OpenStudy (anonymous):

i picked the last one and it was right thank you :)

OpenStudy (jdoe0001):

yw

OpenStudy (whpalmer4):

the last one is the only one that describes the actual roots, so yes, it should be right :-) \[x = -5, x = \pm i\] are the roots \((i = \sqrt{-1})\)

OpenStudy (loser66):

btw, i think we should delete our conversation; it's inappropriate when we occupy other's post to chat. Moreover, ours should be private instead of being here, right?